OpenStudy (anonymous):

Increasing life expectancy. A smoker is 15 times more likely to die of lung cancer than a nonsmoker. An ex-smoker who stopped smoking t years ago is w times more likely to die of lung cancer than a nonsmoker, where w = 15 – t. Graph the equation and use the graph to estimate how much more likely it is for Sandy to die of lung cancer than Polly, if Polly never smoked and Sandy quit 2 1/2 yrs ago.

OpenStudy (anonymous):

polly is nonsmoker sandy is ex- smoker t= 2 1/2= 5/2 years w= 15-t = 15- 5/2 = 25/2 so sandy is 25/2 more times likely to die than polly how would I graph this??? HELP PLEASE!!!!!

OpenStudy (anonymous):

SOMEONE PLEASE HELP ME!!! PLEASE., LAST QUESTION :(